8. A project has an initial cost of $71,950, expected net cash inflows of $9,000 per year for 6 years, and a cost of capital of 10%. What is the project's PI? Do not round your intermediate calculations. Round your answer to two decimal places.

9. A project has an initial cost of $54,925, expected net cash inflows of $12,000 per year for 8 years, and a cost of capital of 11%. What is the project's payback period? Round your answer to two decimal places.

10. A project has an initial cost of $40,000, expected net cash inflows of $9,000 per year for 9 years, and a cost of capital of 11%. What is the project's discounted payback period? Round your answer to two decimal places. Edelman Engineering is considering including two pieces of equipment, a truck and an overhead pulley system, in this year's capital budget. The projects are independent. The cash outlay for the truck is $19,000 and that for the pulley system is $20,000. The firm's cost of capital is 12%. After-tax cash flows, including depreciation, are as follows: Year Truck Pulley 1 $5,100 $7,500 2 5,100 7,500 3 5,100 7,500 4 5,100 7,500 5 5,100 7,500 a. Calculate the IRR for each project. Round your answers to two decimal places. Truck: % What is the correct accept/reject decision for this project? Pulley: % What is the correct accept/reject decision for this project? b. Calculate the NPV for each project. Round your answers to the nearest dollar, if necessary. Enter each answer as a whole number. For example, do not enter 1,000,000 as 1 million. Truck: $ What is the correct accept/reject decision for this project? Pulley: $ What is the correct accept/reject decision for this project? c. Calculate the MIRR for each project. Round your answers to two decimal places. Truck: % What is the correct accept/reject decision for this project? Pulley: % What is the correct accept/reject decision for this project?

11. Talbot Industries is considering launching a new product. The new manufacturing equipment will cost $14 million, and production and sales will require an initial $1 million investment in net operating working capital. The company's tax rate is 35%. a. What is the initial investment outlay? Write out your answer completely. For example, 2 million should be entered as 2,000,000. $ b. The company spent and expensed $150,000 on research related to the new project last year. Would this change your answer? c. Rather than build a new manufacturing facility, the company plans to install the equipment in a building it owns but is not now using. The building could be sold for $1.5 million after taxes and real estate commissions. How would this affect your answer? The project's cost will .

12. The financial staff of Cairn Communications has identified the following information for the first year of the roll-out of its new proposed service: Projected sales $25 million Operating costs (not including depreciation) $12 million Depreciation $6 million Interest expense $3 million The company faces a 40% tax rate. What is the project's operating cash flow for the first year (t = 1)? Write out your answer completely. For example, 2 million should be entered as 2,000,000. $
Answer #1

As per policy, only one question is allowed to answer at a time, here two question are solved :

8) Profitability Index = PV of future inflow / Initial investment
Formula = cash flow per year * PVIFA(cost% , n)
PV of future inflow = $9000 * PVIFA( 10%,6) = 9000 * 4.355 = $39195
Initial investment = $71950
PI of Project = 39195 / 71950 = 0.54
9) Payback period = period in which future inflow covers initial investments
Payback period = Initial investment / Cash inflow per year = 54925/12000 = 4.58 years